# LNS2+RL: Combining Multi-agent Reinforcement Learning with Large Neighborhood Search in Multi-agent Path Finding

### v1

@@ 6.1 Results for multiple map types
LNS2+RL 852.7 942.6 994.9 1052.8
LNS2 831.0 939.4 972.4 1064.2
LACAM 2395.2 5288.3 5035.1 306438.3
EECBS 825.1 912.7 929.7 1030.0
The structural complexity of maps significantly influences algorithm performance. Despite being trained on random maps without prior exposure to highly structured environments like mazes, rooms, and warehouses, LNS2+RL demonstrates remarkable improvements in SR on these three types of maps. Conversely, LACAM excels on less structured maps, achieving a 100% SR on empty maps, but close to 0% SR on highly structured maps. EECBS consistently ranks as the poorest performing algorithm, with SR often remaining at 0% due to the high difficulty of the selected tasks.
Furthermore, taking the tasks in Table 2 as an example, the SoC for LNS2+RL only averages 1.03% higher than that of LNS2. Despite LACAM displaying higher SR on the random-small map, its SoC notably lags behind. Conversely, EECBS consistently achieves the best SoC but at the expense of obtaining the lowest SR. Moreover, LNS2+RL and LNS2 offer the added benefit of maintaining consistent memory usage, in contrast to LACAM and EECBS, whose memory usage escalates rapidly over time due to the generation of longer paths or larger search frontiers.
6.2 Performance analysis
Table 3: Comparing MARL policy against PP+SIPPS on low-level re-planning tasks of MAPF tasks on random-medium map with 65% agent density. The difference between collisions and CP is that for $n$ collisions involving the same agents, the number of collisions is counted as $n$, but CP is counted as $1$. The percentages in parentheses represent the performance ratio of the solution obtained by the MARL policy compared to that of PP+SIPPS.
Iterations Collisions Collision Pairs Sum of Cost Time
PP+SIPPS MARL PP+SIPPS MARL PP+SIPPS MARL PP+SIPPS MARL
1 406.9 106.9(-73.7%) 109.9 93.9(-14.6%) 439.0 592.5(+35.0%) 0.012 0.153(+1175%)
50 471.5 129.9(-72.5%) 107.9 106.4(-1.2%) 564.1 675.3(+19.7%) 0.019 0.159(+737%)
100 488.3 130.8(-73.2%) 101.5 103.8(+2.3%) 597.2 689.9(+15.5%) 0.024 0.176(+633%)
Figure 4: Number of remaining collision pairs through the iterative re-planning process. The dashed blue line shows the switch from MARL to PP+SIPPS.
@@ 6.2 Performance analysis
Table 3 compares the average performance of the MARL policy and PP+SIPPS in replanning tasks across different numbers of iterations. The MARL policy clearly produces fewer collisions overall, but its advantage in reducing CP is less pronounced, indicating that the MARL policy tends to cause collisions with different agents. With re-planning task difficulty decreasing as iterations increase, the MARL policy’s advantage in reducing CP diminishes. Given that the speed of the MARL policy is nearly seven times slower than that of PP+SIPPS, switching to PP+SIPPS becomes a necessary trade-off between time and solution quality.
Figure 4 illustrates the trend of the number of remaining CP over 2000 iterations in MAPF tasks on random-medium maps with a 65% agent density. As expected, in the early stage, before switching to PP+SIPPS, the number of CP for LNS2+RL decreases at a faster rate. After switching, the gap between LNS2+RL and LNS2 continues to widen, as the high solution quality achieved by the MARL policy reduces the difficulty of subsequent replanning tasks.
6.3 Real world experiment
@@
As shown in Figure 1, we implemented our algorithm on a team of 90 virtual and 10 real robots in a warehouse environment. Since during the planning process, our algorithm assumes the position of all agents to be perfectly known at all times, we use ground truth positions for our virtual robots and use external localization (here, the Optitrack Motion Capture System) to obtain accurate position information for our real robots. However, the planned path may not be executed accurately due to disturbances and control inaccuracies. To eliminate these errors, we implement an Action Dependency Graph (ADG), as proposed by Hönig et al. (2019). An ADG introduces a precedence order for the agents occupying a cell to ensure that errors in executing the planned path do not propagate and disrupt the overall execution plan. From our experiment, we observe that agents can reach their goals quickly without collisions, and errors are eliminated by the ADG, demonstrating the potential of using our method in the real world. More information about our experiment is included in Appendix A.4, and the full video is available in the supplementary material.
7 Conclusion
@@
This paper introduces LNS2+RL, a novel MAPF algorithm that combines MARL with large neighborhood search. It follows the overall framework of LNS2, beginning with a path set that may contain collisions and then iteratively replanning paths for subsets of agents until the path set becomes collision-free. In early iterations, LNS2+RL addresses these challenging replanning tasks with a MARL-based planner, which offers higher solution quality but at a slower pace. As iterations progress, it adaptively switches to a fast, priority-based planner to quickly resolve remaining collisions. Our experiments on high-difficulty tasks across various world sizes, team sizes, and map structures consistently demonstrate the superior performance of LNS2+RL compared to other SOTA MAPF algorithms. The practicality of LNS2+RL is validated in a hybrid warehouse mockup simulation involving 10 real-world and 90 simulated robots. However, the MARL planner we use still relies on paths computed by SIPPS, and its SoC cannot be guaranteed as optimal. Future work will explore how to improve this MARL-based planner.
Appendix A Appendix
A.1 Additional results
A.1.1 Detailed results
@@
The detailed results of Section 6.1 can be found in Table 4.
Table 4: Average performance along MAPF metrics of different algorithms for instances with different world sizes, team sizes, and map structure. For maps with sizes $10\times 10$, $25\times 25$, and $50\times 50$, the time constraints for each task are 100 seconds, 600 seconds, and 5000 seconds, respectively, with test counts of 100, 50, and 20. The success rate is the percentage of tasks fully solved within the planning time constraints (i.e., where all agents reached their goal without collisions). The sum of cost only accounts for successful tasks. Partial runtime data for LACAM is unavailable because some tasks in the test exceeded the maximum memory limit we set (100GB), leading to premature termination. The remaining collision pairs metric represents the number of CP remaining in the overall solution after reaching the time constraint and is only available for LNS2+RL and LNS2. Percentages in parentheses represent the reduced remaining CP ratio of LNS2+RL compared to LNS2. ↑ indicates that “higher is better”, and ↓ “lower is better”. ”-” represents unavailable data. The results of the best-performing algorithms in each task are highlighted in bold.
@@
Methods Random-small: 10 * 10 world size with 17.5% static obstacle rate and 50%, 55%, 60%, 65% agents density
Success Rate ↑ Sum of Cost ↓ Runtime ↓ Remaining Collisions Pairs ↓
LNS2+RL 82% 70% 59% 27% 852.7 942.6 994.9 1052.8 21.6 34.9 49.8 79.9
@@
LNS2 80% 68% 50% 27% 831.0 939.4 972.4 1064.2 20.8 33.9 54.1 77.0 0.60 1.51 3.71 10.01
EECBS 38% 18% 3% 1% 825.1 912.7 929.7 1030.0 63.9 82.8 97.1 99.1 - - - -
LACAM 75% 70% 68% 61% 2395.2 5288.3 5035.1 306438.3 27.0 32.2 34.8 42.4 - - - -
Random-medium: 25 * 25 world size with 17.5% static obstacle rate and 50%, 55%, 60%, 65% agents density
LNS2+RL 72% 30% 2% 0% 16726.3 16727.7 16605.0 - 222.1 486.8 593.5 600.0
@@
LACAM 36% 26% 28% 18% 35070.2 70508.0 66149.3 127656.0 397.4 - - - - - - -
Random-large: 50 * 50 world size with 17.5% static obstacle rate and 50%, 55%, 60%, 65% agents density
LNS2+RL 35% 0% 0% 0% 127514.3 - - - 4308.7 5000.0 5000.0 5000.0
@@
LNS2 30% 0% 0% 0% 130756.5 - - - 4329.2 5000.0 5000.0 5000.0 5.15 137.45 1477.10 4261.85
EECBS 0% 0% 0% 0% - - - - 5000.0 5000.0 5000.0 5000.0 - - - -
LACAM 5% 25% 0% 5% 204820.0 955662.6 - 474872.0 - - - - - - - -
Empty: 25 * 25 world size with 0% static obstacle rate and 72.5%, 75%, 77.5%, 80% agents density
LNS2+RL 100% 84% 6% 0% 20155.2 20132.5 19278.0 - 94.7 333.5 596.6 600.0
@@
LACAM 100% 100% 100% 100% 24834.3 26929.9 29408.2 32298.9 0.006 0.007 0.007 0.008 - - - -
Maze: 25 * 25 world size with 45.76% static obstacle rate and 20%, 22.5%, 25%, 27.5% agents density
LNS2+RL 100% 78% 44% 6% 5724.3 6523.3 7307.6 8698.0 16.4 183.6 416.0 568.8
@@
LACAM 2% 0% 0% 0% 41924.0 - - - 588.8 600.0 600.0 - - - - -
Room: 25 * 25 world size with 19.52% static obstacle rate and 40%, 45%, 50%, 55% agents density
LNS2+RL 100% 98% 12% 0% 35497.9 44779.1 58620.2 - 41.3 154.5 577.9 600.0
@@
Warehouse: 25 * 25 world size with 28.8% static obstacle rate and 42.5%, 45%, 47.5%, 50% agents density
LNS2+RL 72% 42% 4% 0% 11202.9 12240.2 12082.0 - 296.9 458.9 595.6 600.0
@@
LACAM 2% 0% 0% 0% 3061736.0 - - - - - - - - - - -
A.1.2 Results for different time limits and easier tasks
@@
In this section, we use the random-small map and warehouse map as representatives of less-structure maps and high-structure maps, respectively, to evaluate the performance of algorithms under varying runtime limits and task difficulties. Specifically, we conducted the runtime-limit experiments on the random-small map with a 60% agent density and the warehouse map with a 42.5% agent density. Their results are shown in Figure 5(a) and 5(b). These results indicate that: First, the relative performance ranking of the algorithms remains consistent across different runtime limits, as discussed in the main paper. LNS2+RL maintains its superiority over LNS2 in both SR and CP. Second, each algorithm has a distinct runtime limit threshold, beyond which its SR stabilizes. Notably, the thresholds of LNS2+RL and LNS2 exceed those of LACAM and EECBS. For instance, as depicted in Figure 5(a), increasing the runtime limit from 20 to 60 seconds results in an 18% increase in SR for LNS2+RL and a 4% increase for LNS2, while the SRs for LACAM and EECBS remain unchanged.
Subsequently, we maintained a constant runtime limit while progressively reducing task difficulty (agent density). The results, illustrated in Figures 5(c) and 5(d), indicate that: First, EECBS and LACAM perform relatively well on less challenging tasks, but the performance gaps between them and LNS2+RL and LNS2 widen as task difficulty escalates. Second, although in simpler tasks, the less pronounced advantages of the MARL-planner over PP+SIPPS lead to an earlier meeting of the switching condition, thus reducing the proportion of MARL-planner usage. However, LNS2+RL continues to achieve performance similar to that of LNS2.
A.1.3 Impact of different hyperparameters
@@
Table 5: Success rate for different neighborhood sizes
@@
Table 6: Success rate for different $\rho$
@@
Table 7: Success rate for different $t_{l}$
@@
Table 8: Success rate for different $t_{h}$
@@
A.2 Map visualization
Figure 6 illustrates examples of the maps used for testing. In all maps, purple cells represent static obstacles, and yellow cells represent empty areas. The static obstacle layouts of the maze, room, and warehouse maps are fixed, and the obstacle layouts of the three random maps are randomly regenerated in each instance. For every instance of all maps, the agents’ distinct starting and goal positions are randomly chosen from empty cells within the same connected region of the map. The tasks tested in Section 6.1 are challenging, with the vacancy rates for the most difficult tasks on the seven maps only being 17.5%, 17.5%, 17.5%, 20%, 26.74%, 25.48%, and 21.2% respectively.
(a) Random-small
(b) Random-medium
(c) Random-large
(d) Empty
(e) Maze
(f) Room
(g) Warehouse
A.3 Hyperparameters
Table 9: Hyperparameters table
@@ A.3 Hyperparameters
Training timestep for the first stage 7e6
Training timestep for the second stage 7e6
@@
Table 9 shows the hyperparameters used to train the MARL model and implement the LNS2+RL algorithm evaluated in Section 6 of our paper. The first 29 hyperparameters are those used to train the MARL model, while the remaining hyperparameters are used to insert the MARL planner into the overall framework. There are three different task difficulties in the curriculum learning of training stage one, so ratio of total number of agents to world size, obstacle density, $\alpha$, Imitation learning timestep and maximum iterations contain three values/value sets, each value/value set corresponding to a task difficulty. The hyperparameters used in the second stage of training are mostly the same as those in the first stage of training, except ratio of total number of agents to world size = (60%, 65%, 70%), obstacle density = (15%, 17.5%, 20%),$\alpha$ = 0.04,maximum iterations = 100
A.4 Real world experiment
A.4.1 Setup
@@
We use a warehouse environment for our experiment (Figure 7). We assume that each pixel has a resolution of 0.5 m per grid unit, and based on this resolution, we realize a 7m x 4.5m part of the warehouse in the real world. We use a total of 100 robots for our experiment, of which 90 are virtual, and 10 are real Sparklife Omnibots robots with dimensions of approximately 0.3m ×0.3m equipped with mecanum wheels for omnidirectional movement. We randomly generate the start and end positions of robots in the environment, with the constraint that the start and end positions of the real robots are in the real part of the warehouse. Since the MAPF problem assumes the accurate position of all agents for all time, we considered the ground truth positions for all virtual robots and used the OptiTrack Motion Capture System to get accurate positions for our real robots. Finally, we pre-plan the complete path set using LNS2+RL before actual execution and compute the joint actions agents should take at each timestep based on this path set.
A.4.2 Action dependency graph
The paths generated by LNS2+RL assume perfect operation for each agent at every step in the discrete map. However, due to the robots being imperfect and the environment being continuous, executing the planned path directly in the real world is infeasible.
Figure 7: Map of the 23 x 23 warehouse environment used in our experiments, where black cells represent obstacles and white cells free space.
@@ A.4.2 Action dependency graph
For instance, a planner may output a set of actions that lets an agent (say ’A’) take another agent’s (say ’B) position while Agent B moves to a new cell. If we were to try and execute this behavior on robots directly, a number of things, like errors in localization, delays in executing motor control commands, or differences in velocities, could lead to a collision between agents. To avoid such scenarios and make our joint set of actions executable, we follow the method proposed by Hönig et al. (2019) and construct an Action Dependency Graph (ADG). The main benefit of having an ADG is that it introduces a precedence order for the agents occupying a cell. This means that if an agent moves faster than the remaining agents, it will still wait for agents to catch up if it occupies a cell after another agent in the originally planned path. This behavior ensures that the errors in executing the planned path do not propagate and hamper the execution plan. Returning to our earlier example, using an ADG ensures that since Agent A is headed to a cell that Agent B occupies before it is in the originally planned path, it will wait for Agent B to reach a new cell before moving to occupy it.
A.4.3 Execution

### v2

@@ 4. Alleviate congestion
50% agent
density
55% agent
density
60% agent
density
65% agent
density
LNS2+RL 852.7 942.6 994.9 1,052.8
LNS2 831.0 939.4 972.4 1,064.2
LaCAM 2,395.2 5,288.3 5,035.1 306,438.3
EECBS 825.1 912.7 929.7 1,030.0
SCRIMP 3,719.4 4,444.3 9,869.7 -
Table 2: Sum of costs on random-small map (accounting only for successful tasks). ”-” represents unavailable data due to 0% SR.
Performance Analysis
@@
Table 3 shows the performance ratio of the solution obtained by the MARL policy compared to that of PP+SIPPS in replanning tasks across different numbers of iterations. The MARL policy clearly produces fewer collisions overall, but its advantage in reducing CP is less pronounced, indicating that the MARL policy tends to cause collisions with different agents. With replanning task difficulty decreasing as iterations increase, the MARL policy’s advantage in reducing CP diminishes. Given that the speed of the MARL policy is nearly seven times slower than that of PP+SIPPS, switching to PP+SIPPS becomes a necessary trade-off between time and solution quality (the number of collisions in the solution). Figure 3 illustrates the trend of the number of remaining CP over 2000 iterations in MAPF tasks on random-medium maps with a 65% agent density. As expected, in the early stage, before switching to PP+SIPPS, the number of CP for LNS2+RL decreases at a faster rate. After switching, the gap between LNS2+RL and LNS2 continues to widen, as the high solution quality achieved by the MARL policy reduces the difficulty of subsequent replanning tasks.
Iterations Collisions Colliding Pairs Sum of Cost Time
1 -73.7% -14.6% +35.0% +1,175%
50 -72.5% -1.2% +19.7% +737%
100 -73.2% +2.3% +15.5% +633%
Table 3: Performance ratio of the solution obtained by the MARL policy compared to that of PP+SIPPS on low-level replanning tasks in random-medium maps with 65% agent density. The difference between collisions and CP is that for $n$ collisions involving the same agents, the number of collisions is counted as $n$, but CP is counted as $1$.
Figure 3: Number of remaining CP through the iterations. The dashed blue line shows the switch from MARL to PP+SIPPS.
Real World Experiment
@@
We implemented our algorithm on a team of 90 virtual and 10 real robots in a warehouse environment. The planned path may not be executed accurately due to disturbances and control inaccuracies. To eliminate these errors, we implement an Action Dependency Graph (ADG) (Hönig et al. 2019). From our experiment, we observe that agents can reach their goals quickly without collisions, and errors are eliminated by the ADG, demonstrating the potential of using our method in the real world. More information about our experiment are included in the supplementary material.
Conclusion
@@
This paper introduces LNS2+RL, a novel MAPF algorithm that combines MARL with large neighborhood search. It follows the overall framework of LNS2, beginning with a path set that may contain collisions and then iteratively replanning paths for subsets of agents until the path set becomes collision-free. In early iterations, LNS2+RL addresses these challenging replanning tasks with a MARL-based planner, which offers higher solution quality (fewer collisions) but at a slower pace. As iterations progress, it adaptively switches to a fast, prioritized planning algorithm to quickly resolve remaining collisions. Our experiments on high-agent-density tasks across various world sizes, team sizes, and map structures consistently demonstrate the superior performance of LNS2+RL compared to other state-of-the-art MAPF algorithms. The practicality of LNS2+RL is validated in a hybrid warehouse mockup simulation.
Acknowledgments
This research was supported by the Singapore Ministry of Education (MOE), as well as by an Amazon Research Award.
Additional Results
Detailed Results
@@ Acknowledgments
The detailed results of the Experiment Section in the main body of the paper can be found in Table 4.
@@
Methods Random-small: 10 * 10 world size with 17.5% static obstacle rate and 50%(50 agents), 55%(55 agents), 60%(60 agents), 65%(65 agents) agents density
Success Rate ↑ Sum of Cost ↓ Runtime ↓ Remaining Colliding Pairs ↓
LNS2+RL 82% 70% 59% 27% 852.7 942.6 994.9 1,052.8 21.6 34.9 49.8 79.9
@@
LNS2 80% 68% 50% 27% 831.0 939.4 972.4 1,064.2 20.8 33.9 54.1 77.0 0.60 1.51 3.71 10.01
EECBS 38% 18% 3% 1% 825.1 912.7 929.7 1,030.0 63.9 82.8 97.1 99.1 - - - -
LaCAM 75% 70% 68% 61% 2,395.2 5,288.3 5,035.1 306,438.3 27.0 32.2 34.8 42.4 - - - -
SCRIMP 22% 12% 3% 0% 3,719.4 4,444.3 9,869.7 - 2.7 3.3 4.2 5.1 - - - -
Random-medium: 25 * 25 world size with 17.5% static obstacle rate and 50%(312 agents), 55%(344 agents), 60%(375 agents), 65%(406 agents) agents density
LNS2+RL 72% 30% 2% 0% 16,726.3 16,727.7 16,605.0 - 222.1 486.8 593.5 600.0
@@
LaCAM 36% 26% 28% 18% 35,070.2 70,508.0 66,149.3 127,656.0 397.4 - - - - - - -
SCRIMP 0% 0% 0% 0% - - - - 96.8 140.9 210.2 284.5 - - - -
Random-large: 50 * 50 world size with 17.5% static obstacle rate and 50%(1,250 agents), 55%(1,375 agents), 60%(1,500 agents), 65%(1,625 agents) agents density
LNS2+RL 35% 0% 0% 0% 127,514.3 - - - 4,308.7 5,000.0 5,000.0 5,000.0
@@
LNS2 30% 0% 0% 0% 130,756.5 - - - 4,329.2 5,000.0 5,000.0 5,000.0 5.15 137.45 1,477.10 4,261.85
EECBS 0% 0% 0% 0% - - - - 5,000.0 5,000.0 5,000.0 5,$0$
LaCAM 5% 25% 0% 5% 204,820.0 955,662.6 - 474,872.0 - - - - - - - -
SCRIMP 0% 0% 0% 0% - - - - - - - - - - - -
Empty: 25 * 25 world size with 0% static obstacle rate and 72.5%(453 agents), 75%(469 agents), 77.5%(484 agents), 80%(500 agents) agents density
LNS2+RL 100% 84% 6% 0% 20,155.2 20,132.5 19,278.0 - 94.7 333.5 596.6 600.0
@@
LaCAM 100% 100% 100% 100% 24,834.3 26,929.9 29,408.2 32,298.9 0.006 0.007 0.007 0.008 - - - -
SCRIMP 0% 0% 0% 0% - - - - 167.8 217.3 286.4 346.5 - - - -
Maze: 25 * 25 world size with 45.76% static obstacle rate and 20%(125 agents), 22.5%(141 agents), 25%(156 agents), 27.5%(172 agents) agents density
LNS2+RL 100% 78% 44% 6% 5,724.3 6,523.3 7,307.6 8,698.0 16.4 183.6 416.0 568.8
@@
LaCAM 2% 0% 0% 0% 41,924.0 - - - 588.8 600.0 600.0 - - - - -
SCRIMP 0% 0% 0% 0% - - - - 14.1 19.2 25.2 33.6 - - - -
Room: 25 * 25 world size with 19.52% static obstacle rate and 40%(250 agents), 45%(281 agents), 50%(321 agents), 55%(344 agents) agents density
LNS2+RL 100% 98% 12% 0% 35,497.9 44,779.1 58,620.2 - 41.3 154.5 577.9 600.0
@@
SCRIMP 0% 0% 0% 0% - - - - 85.8 123.4 157.1 200.9 - - - -
Warehouse: 25 * 25 world size with 28.8% static obstacle rate and 42.5%(266 agents), 45%(281 agents), 47.5%(297 agents), 50%(312 agents) agents density
LNS2+RL 72% 42% 4% 0% 11,202.9 12,240.2 12,082.0 - 296.9 458.9 595.6 600.0
@@
LaCAM 2% 0% 0% 0% 3,061,736.0 - - - - - - - - - - -
SCRIMP 0% 0% 0% 0% - - - - 60.8 73.2 92.3 114.6 - - - -
Table 4: Average performance along MAPF metrics of different algorithms for instances with different world sizes, team sizes, and map structure. For maps with sizes $10\times 10$, $25\times 25$, and $50\times 50$, the test counts are set to 100, 50, and 20, respectively, and the time constraints for each task, except when using SCRIMP, are set to 100 seconds, 600 seconds, and 5,000 seconds. To ensure consistency with the original paper, SCRIMP uses the maximum timestep as the termination condition, with maximum timesteps of 356, 556, and 1024 for maps sized $10\times 10$, $25\times 25$, and $50\times 50$, respectively. The success rate is the percentage of tasks fully solved before the termination condition is met(i.e., where all agents reached their goal without collisions). The sum of cost only accounts for successful tasks. Partial runtime data for LaCAM and SCRIMP is unavailable because some tasks in the test exceeded the maximum memory limit we set (100GB), leading to premature termination. The remaining colliding pairs metric represents the number of CP remaining in the overall solution after reaching the time constraint and is only available for LNS2+RL and LNS2. Percentages in parentheses represent the reduced remaining CP ratio of LNS2+RL compared to LNS2. ↑ indicates that “higher is better”, and ↓ “lower is better”. ”-” represents unavailable data. The best-performing algorithms in each task are highlighted in bold. Because SCRIMP uses maximum timestep as the termination condition, it is excluded from the selection of the shortest time.
Results for Different Time Limits and Easier Tasks
@@
In this section, we use the random-small map and warehouse map as representatives of maps with simple obstacle structure and maps with complex obstacle structure, respectively, to evaluate the performance of algorithms under varying runtime limits and task difficulties. SCRIMP is not included in the runtime-limit experiment due to its different termination condition. Specifically, we conducted the runtime-limit experiments on the random-small map with a 60% agent density and the warehouse map with a 42.5% agent density. Their results are shown in Figure 4(a) and 4(b). These results indicate that: First, the relative performance ranking of the algorithms remains consistent across different runtime limits, as discussed in the main paper. LNS2+RL maintains its superiority over LNS2 in both SR and CP. Second, each algorithm has a distinct runtime limit threshold, beyond which its SR stabilizes. Notably, the thresholds of LNS2+RL and LNS2 exceed those of LaCAM and EECBS. For instance, as depicted in Figure 4(a), increasing the runtime limit from 20 to 60 seconds results in an 18% increase in SR for LNS2+RL and a 4% increase for LNS2, while the SRs for LaCAM and EECBS remain unchanged.
Subsequently, we maintained a constant runtime limit/maximum timestep for SCRIMP while progressively reducing task difficulty (agent density). The results, illustrated in Figures 4(c) and 4(d), indicate that: First, EECBS and LaCAM perform relatively well on less challenging tasks, but the performance gaps between them and LNS2+RL and LNS2 widen as task difficulty escalates. SHRIMP still performs the worst across all task difficulties. Second, although in simpler tasks, the less pronounced advantages of the MARL-planner over PP+SIPPS lead to an earlier meeting of the switching condition, thus reducing the proportion of MARL-planner usage. However, LNS2+RL continues to achieve performance similar to that of LNS2.
Impact of Different Hyperparameters
@@
Table 5: Success rate for different neighborhood sizes
@@
Table 6: Success rate for different $\rho$
@@
Table 7: Success rate for different $t_{l}$
@@
Table 8: Success rate for different $t_{h}$
The Importance of Using SIPPS+PP for Path Supplementation
Iterations Number of Agents Reached Goal One-Shot Success Rate
$t_{l}$=0.8 $t_{l}$=1 $t_{l}$=1.2 $t_{l}$=1.4 $t_{l}$=0.8 $t_{l}$=1 $t_{l}$=1.2 $t_{l}$=1.4
1 6.12 7.12 7.50 7.59 25% 64% 78% 80%
50 6.01 6.96 7.37 7.47 22% 51% 69% 76%
100 6.01 6.89 7.31 7.48 23% 51% 69% 76%
Table 9: Goal-reaching agents vary with number of iterations and $t_{l}$
Since the MARL-based planner cannot guarantee solution completeness, LNS2+RL employs SIPPS+PP to supplement the remaining paths for agents that have not reached their goals by timestep $t_{l}$, thereby fully leveraging the MARL-planner’s capabilities. In this section, we present how the PMDO task one-shot success rate of our MARL-based planner varies with the number of iterations and $t_{l}$ on random-small maps with a 60% agent ratio and a neighborhood size of 8, to demonstrate the importance of using SIPPS+PP for path supplementation. As shown in Table 9, the success rate (the proportion of PMDO tasks where all agents reach their goals) of our MARL-based planner at $t_{l}=1.2$ and the first iteration is 78%. As the number of iterations increases, the task difficulty gradually rises, leading to a decline in the success rate. The number of agents reaching their goal per task remains around 7.5, indicating that the task failures are caused by a small number of agents failing to find paths. This observation suggests that updating the overall path set solely with solutions where all agents reach their goals could result in the loss of many high-quality solutions already obtained. Additionally, we note that the benefits of increasing $t_{l}$ diminish over time. For example, at the 50th iteration, increasing $t_{l}$ from 1 to 1.2 raises the average number of goal-reaching agents by 0.41, whereas increasing it from 1.2 to 1.4 only yields an additional 0.1 agents. Therefore, we set $t_{l}=1.2$ to balance planning time and gains.
Number of Iterations
Algorithms 50 agents 55 agents 60 agents 65 agents
LNS2+RL 364,549.8 377,825.4 410,457.8 407,344.2
LNS2 373,025.6 382,657.5 451,732.3 453,540.0
Table 10: Average number of iterations on random-small maps with different numbers of agents.
Table 10 shows the average number of iterations spent by LNS2+RL and LNS on random-small maps with different numbers of agents. The results indicate that, compared to LNS, LNS2+RL requires fewer iterations to solve MAPF tasks.
Hyperparameters
@@
Training timestep for the first stage 7e7
Training timestep for the second stage 7e7
@@
Table 11 shows the hyperparameters used to train the MARL model and implement the LNS2+RL algorithm evaluated in Experiment Section of our paper. The last four hyperparameters are those used to insert the MARL planner into the overall framework, while the remaining hyperparameters are used to train the MARL model. There are three different task difficulties in the curriculum learning of training stage one, so ratio of total number of agents to world size, obstacle density, $\alpha$, Imitation learning timestep and maximum iterations contain three values/value sets, each value/value set corresponding to a task difficulty. The hyperparameters used in the second stage of training are mostly the same as those in the first stage of training, except ratio of total number of agents to world size = (60%, 65%, 70%), obstacle density = (15%, 17.5%, 20%),$\alpha$ = 0.04,maximum iterations = 100
Training Details
The training was conducted on a server equipped with four Nvidia GeForce RTX 3090 GPUs and one Intel(R) Core(TM) i9-10980XE CPU (18 cores, 36 threads). Only one GPU was used during training. The code for the neural network part was written in PyTorch 2.1.1. We relied on Ray 3.00 to employ 32 processes to collect data in parallel. Both training stages one and two lasted 7e7 timesteps, each requiring 136 hours. However, the training curve in stage one converged at 4e7 timesteps, and in stage two at 5e7 timesteps. We extended the training duration to achieve a model with optimal performance, which can scale to different numbers of agents and adapt to different map structures without retraining.
Summary of Observation Channels
As an example at timestep $t$, Table 12 shows the information that agent $i$ will observe.
No. Information included
1 Positions of hard obstacles.
2 Positions of other agents.
3 Actual/projected positions of observable agents’ goals.
4 Position of agent $i$’s own goal.
5, 6, 7, 8
There are four channels in total, each representing a move action (up, down, left, right).
When considering only hard obstacles, the value is 1 if selecting the corresponding
action at a given position brings agent $i$ closer to its goal; otherwise, it is 0.
9 SIPPS map. All cells traversed by agent $i$’s SIPPS path $\hat{p}_{i}$ during $t-15$ to $t+15$.
10
Occupancy duration map. The ratio of time a cell will be occupied by obstacles
to the predefined episode termination length.
11
Blank duration map. The ratio of time a cell will remain obstacle-free
to the predefined episode termination length.
12, 13, 14, 15, 16, 17, 18, 19, 20 Future path maps of soft obstacles. The positions of soft obstacles from $t$ to $t+8$.
21, 22, 23, 24, 25 Predicted path maps of other agents. The predicted positions of other agents from $t+1$ to $t+5$.
26
Cell utilization map. The normalized sum of the number of timesteps soft obstacles
occupy the cell from $t-2$ to $t+15$, and the number of timesteps agents occupy the cell
from $t-2$ to $t$.
27, 28, 29, 30, 31
Direction utilization maps. There are five channels in total, each representing an action
(up, down, left, right, stay). The normalized sum of the number of timesteps soft obstacles
traverse the cell and select the corresponding action from timestep $t-2$ to $t+15$, and
the number of timesteps agents traverse the cell and select the corresponding action
from timestep $t-2$ to $t$.
Table 12: The observation channels of agent $i$ at time step $t$ and their corresponding information.
Map Visualization
Figure 5 illustrates examples of the maps used for testing. In all maps, purple cells represent static obstacles, and yellow cells represent empty areas. The static obstacle layouts of the maze, room, and warehouse maps are fixed, and the obstacle layouts of the three random maps are randomly regenerated in each instance. For every instance of all maps, the agents’ distinct starting and goal positions are randomly chosen from empty cells within the same connected region of the map. The tasks tested in Experiment Section are challenging, with the vacancy rates for the most difficult tasks on the seven maps only being 17.5%, 17.5%, 17.5%, 20%, 26.74%, 25.48%, and 21.2% respectively.
(a) Random-small
(b) Random-medium
(c) Random-large
(d) Empty
(e) Maze
(f) Room
(g) Warehouse
Real World Experiment
Figure 6: Real portion of the warehouse with 10 robots (left), and complete rviz visualization with 100 robots (right). Real robots are circled in different colors, and crosses of the same colors represent their goal positions
Setup
@@
We implemented our algorithm on a team of 90 virtual and 10 real robots in a warehouse environment (Figure 6). We assume that each pixel has a resolution of 0.5 m per grid unit, and based on this resolution, we realize a 7m x 4.5m part of the warehouse in the real world. The 10 real robots are Sparklife Omnibots robots with dimensions of approximately 0.3m ×0.3m equipped with mecanum wheels for omnidirectional movement. We randomly generate the start and end positions of robots in the environment, with the constraint that the start and end positions of the real robots are in the real part of the warehouse Since the MAPF problem assumes the accurate position of all agents for all time, we considered the ground truth positions for all virtual robots and used the OptiTrack Motion Capture System to get accurate positions for our real robots. Finally, we pre-plan the complete path set using LNS2+RL before actual execution and compute the joint actions agents should take at each timestep based on this path set.
Action Dependency Graph
@@
The paths generated by LNS2+RL assume perfect operation for each agent at every step in the discrete map. However, due to the robots being imperfect and the environment being continuous, executing the planned path directly in the real world is infeasible. For instance, a planner may output a set of actions that lets an agent (say ’A’) take another agent’s (say ’B) position while Agent B moves to a new cell. If we were to try and execute this behavior on robots directly, a number of things, like errors in localization, delays in executing motor control commands, or differences in velocities, could lead to a collision between agents. To avoid such scenarios and make our joint set of actions executable, we follow the method proposed by (Hönig et al. 2019) and construct an Action Dependency Graph (ADG). The main benefit of having an ADG is that it introduces a precedence order for the agents occupying a cell. This means that if an agent moves faster than the remaining agents, it will still wait for agents to catch up if it occupies a cell after another agent in the originally planned path. This behavior ensures that the errors in executing the planned path do not propagate and hamper the execution plan. Returning to our earlier example, using an ADG ensures that since Agent A is headed to a cell that Agent B occupies before it is in the originally planned path, it will wait for Agent B to reach a new cell before moving to occupy it.
Execution